PLEASE HELP.
A student makes a $9.50 purchase at the bookstore with a $20 bill. The store has no bills and given the change in quarters and fifty cent prices. There are 30 coins in all. How many of each kind are there?

Answer by Jk22(389)

let f be the number of fifty cents coins, then there are (30-f) quarter, we shall have :
f*0.5+(30-f)*.25 = (20-9.5) = 10.5$

f/2 + 30/4 - f/4 = 10.5

f/4 = 21/2 - 30/4 = (42 - 30)/4

f = 12 coins of fifty cents and 18 coins of a quarter $.